OpenStudy (anonymous):

Evaluate when x = 9.

OpenStudy (isaiah.feynman):

Evaluate what?

OpenStudy (anonymous):

\[\sqrt{x+7} - \sqrt{3x}\]

OpenStudy (anonymous):

sqrt(16)-sqrt(27) =4 - 3sqrt(3)

OpenStudy (isaiah.feynman):